How would the use of an economic pricing strategy (profit-maximizing rule or price discrimination) improve the operational profitability of most organizations in the current economic environment? How would the practice benefit the business or organization for which you are working? Explain.

no less than 250 words in length, make at least one reference to your text or other course materials and provide in-text citations. As you reference information from a source, be sure to provide APA citations in text and at the end of your post.
Answer #1

Answer :

Economic pricing strategy is the policy which makes it possible for the firm to earn the maximum economic profit. If a price is higher than average cost (AC), then economic profit is earned. Here AC includes the cost of four factors used in production. They are Rent for using land, wages for labor service, interest for capital and normal profit for an organizer. When AC and AR are equal, only normal profit is earned.

A primary objective of a firm is to maximize profit. Any quantity of production cannot help the firm to achieve this objective. It requires the production of a specific quantity. The condition of profit maximization is-

rule 1: the First derivative of profit function with respect to quantity is zero. Here profit is the excess of total revenue(TR) over total cost (TC)

\textup{Profit} (\pi )=\textup{TR}-\textup{TC}

\therefore \frac{d\pi }{dQ}=\frac{d}{dQ}\left[\textup{TR}\right]-\frac{d}{dQ}\left[\textup{TC}\right]=0

Here the first derivative of the total function is marginal value. It indicates the change in TR or TC due to a unit change in quantity.

\therefore \frac{d\pi }{dQ}=MR-MC=0

\therefore MR=MC.......................................................(1)

This condition requires that MR and MC should be equal. Suppose the cost of producing one extra unit is $6. It includes the cost of all factors used in this extra product. Normal profit is also included here as a cost of an organizer. If this extra unit can be sold at $10, then MR is $10. So by selling this extra unit, a firm will be able to add $4 profit. In this situation, a firm will go on producing more and more units. Ultimately it will stop production where MC and MR are equal. If the firm produces after this, then will exceed MR. So total profit will decline. Diagrammatically, this condition is satisfied, when MC curve intersects MR curve.

Condition 2:

Satisfaction of the first condition will mean either profit has been maximized or minimized. So for getting confirmed result, you have to satisfy sufficient condition. It requires that second derivative of profit function should be negative. Thus-

\frac{d}{dQ}\left[MR-MC\right]<0

\therefore \frac{d}{dQ}\left[MR\right]-\frac{d}{dQ}\left[MC \right ]<0

\therefore \frac{d}{dQ}\left[MR\right]<\frac{d}{dQ}\left[MC \right ]

Here left side indicates the slope of MR curve and the right side is a slope of MC curve. This the second condition requires that MC curve must cut MR from the below. If this condition is fulfilled, then slope of MCmustbe greater than the slope of MR.

Thus fulfilment of these two conditions will help the firm to improve the operational profitability.

Sometimes, it is possible to distinguish a market or its customers in two or more groups. Elasticity of demand of such groups are different. In that situation, the firm can chargethe different price to different customers. Customers with Inelastic demand can be charged the higher price. Due to inelasticity, demand will fall less thanan increase in price. So total profit will increase. For other group, pthe rice will be reduced. Then demand will increase more thanthe price cut. Result will be ria se in profit again. This pricing strategy is known as price discrimination. It also helps the firm in improving profitability.

The practice of profit maximization will benefit the organization. Modern concept is that a firm should maximize value. It is possible if it can earn something more than normal profit. When AR and AC are equal, then only normal profit is earned . It is the minimum required the profit for survival. So anything earned above normal profit will add value to the firm. Economic pricing strategy will help the firm to achieve it.